### v1

# Does SuperKamiokande Observe Lévy Flights of Solar Neutrinos?

Abstract. The paper is utilyzing data from the SuperKamiokande solar neutrino detection experiment and analyses them by diffusion entropy analysis and standard deviation analysis. The result indicates that solar neutrinos are subject to Lévy flights (anomalous diffusion, superdiffusion). Subsequently the paper derives the probability density function and the governing fractional diffusion equation (fractional Fokker-Planck Equation) for solar neutrino Lévy flights.
Keywords: solar neutrinos, SuperKamiokande experiment data, diffusion entropy analysis, standard deviation analysis, superdiffusion, scaling exponent, neutrino probability density function, Fox H-function, fractional diffusion equation.
1 Solar Neutrinos: SuperKamiokande Data
Over the past 50 years, radio-chemical and real-time solar neutrino experiments have proven to be sensitive tools to test both astrophysical and elementary particle physics models and principles (Sakurai, 2018; Orebi Gann et al., 2021). Solar neutrino detectors (radio-chemical: Homestake, GALLEX + GNO, SAGE, real-time: SuperKamiokande, SNO, Borexino) have demonstrated that the Sun is powered by thermonuclear fusion reactions. Two distinct processes, the pp-chain and the sub-dominant CNO-cycle, are producing solar neutrinos with different energy spectra and fluxes (see Figure 1). To date fluxes from the pp-chain have been measured: ${}^{7}Be$, ${}^{8}B$, and, indirectly, $pp$. Experiments with solar neutrinos and reactor anti-neutrinos (KamLAND) have confirmed that solar neutrinos undergo flavor oscillations (Mikheyev-Smirnov-Wolfenstein (MSW) model). Results from solar neutrino experiments are consistent with the Mikheyev-Smirnov-Wolfenstein Large Mixing Angle (MSW-LMA) model, which predicts a transition from vacuum-dominated to matter-enhanced oscillations, resulting in an energy dependent electron neutrino survival probability. Non-standard neutrino interaction models derived such neutrino survival probability curves that deviate significantly from MSW-LMA, particularly in the 1-4 $MeV$ transition region.
Figure 1: The solar neutrino spectrum for the pp-chain and the CNO-cycle and parts of the spectrum that are being detectable by the experiments based on gallium, chlorine, and Cherenkov radiation
Figure 2: The variation of the solar neutrino flux over time as shown in the SuperKamiokande I, II, and III experiments
2 Diffusion Entropy and Standard Deviation: Analysis
For all radio-chemical and real-time solar neutrino experiments, periodic variation in the detected solar neutrino fluxes have been reported, based mainly on Fourier and wavelet analysis methods (standard deviation analysis). Other attempts to analyze the same data sets, particularly undertaken by the experimental collaborations of real-time solar neutrino experiments themselves, have failed to find evidence for such variations of the solar neutrino flux over time (Abe et al., 2023). Periodicities in the solar neutrino fluxes, if confirmed, could provide evidence for new solar, nuclear, or neutrino physics beyond the commonly accepted physics of vacuum-dominated and matter-enhanced oscillations of massive neutrinos (MSW model) that is, after 50 years of solar neutrino experiment and theory, considered to be the ultimate solution to the solar neutrino problem. Specifically, subsequent to the analysis made by the SuperKamiokande collaboration, the SNO experiment collaboration has painstakingly searched for evidence of time variability at periods ranging from 10 $years$ down to 10 $minutes$. SNO has found no indications for any time variability of the ${}^{8}B$ flux at any timescale, including in the frequency window in which $g$-mode oscillations of the solar core might be expected to occur. Despite large efforts to utilize helio-seismology and helio-neutrinospectroscopy, at present time there is no conclusive evidence in terms of physics for time variability of the solar neutrino fluxes from any solar neutrino experiment. If such a variability over time would be discovered, a mechanism for a chronometer for solar variability could be proposed based on relations between properties of thermonuclear fusion and g-modes (Buldgen et al., 2024; Sturrock et al., 2021). All above findings encouraged the conclusion that Fourier and wavelet analysis, which are based upon the analysis of the variance of the respective time series (standard deviation analysis: SDA) should be complemented by the utilization of diffusion entropy analysis (DEA), which measures the scaling of the probability density function (pdf) of the diffusion process generated by the time series thought of as the physical source of fluctuations (Scafetta, 2010). For this analysis, we have used the publicly available data of SuperKamiokande-I and SuperKamiokande-II (see Figure 2) (Yoo et al., 2003: Cravens et al., 2008; Abe et al., 2024). Such an analysis does not reveal periodic variations of the solar neutrino fluxes but shows how the pdf scaling exponent departs in the non-Gaussian case from the Hurst exponent. Figures 3 to 6 show the Hurst exponents (SDA) and scaling exponents (DEA) for the SuperKamiokande I and II data. SuperKamiokande is sensitive mostly to neutrinos from the ${}^{8}B$ branch of the $pp$ nuclear fusion chain in solar burning. Above approximately 4 $MeV$ the detector can pick-out the scattering of solar neutrinos off atomic electrons which produces Cherenkov radiation in the detector. The ${}^{8}B$ and rarer $hep$ neutrinos have a spectrum which ends near 20 $MeV$ (see Figure 1).
Assuming that the solar neutrino signal is governed by a probability density function with scaling given by the asymptotic time evolution of a pdf of $x$, obeying the property (Scafetta, 2010; Culbreth et al, 2023)
$p(x,t)=\frac{1}{t^{\delta}}f(\frac{x}{t^{\delta}}),$ (1)
where $\delta$ denotes the scaling exponent of the pdf. In the variance based methods, scaling is studied by direct evaluation of the time behavior of the variance of the diffusion process. If the variance scales, one would have
$\sigma_{x}^{2}(t)\sim t^{2H},$ (2)
where $\sigma_{x}^{2}(t)$ is the variance of the diffusion process and where $H$ is the Hurst exponent. To evaluate the Shannon entropy of the diffusion process at time $t$, defined $S(t)$ as
[…]
and with the previous $p(x,t)$ one has
[…]
The scaling exponent $\delta$ is the slope of the entropy against the logarithmic time scale. The slope is visible in Figures 4 and 6 for the SuperKamiokande data measured for ${}^{8}B$ and $hep$. The Hurst exponents (SDA) are $H=0.66$ and $H=0.36$ for ${}^{8}B$ and $hep$, respectively, shown in Figures 3 and 5. The $pdf$ scaling exponents (DEA) are $\delta=0.88$ and $\delta=0.80$ for ${}^{8}B$ and $hep$, respectively, as shown in Figures 4 and 6. The values for both SDA and DEA indicate a deviation from Gaussian behavior which would require that $H=\delta=0.5$. A test computation for the application of SDA and DEA to data that are known to exhibit non-Gaussian behavior have been published by Haubold et al. (2012) and Tsallis (2024).
In this test, SDA and DEA, applied to the magnetic field strength fluctuations recorded by the Voyager-I spacecraft in the heliosphere clearly revealed the scaling behavior of such fluctuations as previously already discovered by non-extensive statistical mechanics considerations that lead to the determination of the non-extensivity q-triplet
3 Probability Density Function and Differential Equation: Lévy Flights
We consider a diffusion process generated by a waiting time pdf with a finite characteristic time $T$ that can be modeled with $a$ Poissonian distribution, and a jump length pdf $\lambda(x)$ given by a Lévy distribution with index $20<\alpha<$ 2 (Metzler and Klafter, 2000). The Fourier transform of $\lambda(x)$ is
$\hat{\lambda}(k)=exp(-\sigma^{\alpha}|k|^{\alpha})\sim 1-\sigma^{\alpha}|k|^{\alpha}.$ (5)
Then $\lambda(x)$ has the asymptotic behavior given by
$\lambda(x)\sim A_{\alpha}\sigma^{\alpha}|x|^{-1-\alpha}=A_{\alpha}\sigma^{1-\mu}|x|^{-\mu}$ (6)
for $|x|\gg\sigma$ and $\mu=1+\alpha$. Substituting the asymptotic expansion of the jump length pdf $\hat{\lambda}(k)$ in the Fourier space and the waiting time pdf of
$\Phi(t)=\frac{1}{\tau}exp(-\frac{t}{\tau})$ (7)
where $\tau=T<\infty$ is the characteristic waiting time in the Laplace space into
$\hat{p}(k,s)=\frac{1-\hat{\Phi}(s)}{s}\frac{\hat{p_{0}}(k)}{1-\hat{\Phi}(s)\hat{\lambda}(k)},$ (8)
where $\hat{p_{0}}(k)$ is the Fourier transform of the initial condition $p(x,0)$, we obtain the following jump pdf in the Fourier-Laplace space
$\hat{p}(k,s)=\frac{1}{s+K^{\alpha}|k|^{\alpha}},$ (9)
where $K^{\alpha}=\sigma^{\alpha}/\tau$ is the generalized diffusion constant. Eq.(9) is the solution of the generalized diffusion equation (Hilfer, 2018)
$\frac{\partial p(x,t)}{\partial t}=K^{\alpha}\;_{-\infty}D^{\alpha}_{x}p(x,t),$ (10)
where ${}_{-\infty}D^{\alpha}_{x}$ is the fractional Weyl operator.
Upon Laplace inversion of Eq. (9), we get the characteristic function of the jump pdf
$\hat{p}(k,t)=exp(-K^{\alpha}t|k|^{\alpha}).$ (11)
Eq.(11) is the characteristic function of a centered and symmetric Lévy distribution. The Fourier inversion of (11) can be analytically obtained by making use of the Fox function (Mathai et al., 2010; Mathai and Haubold, 2018)
$p(x,t)$ $=\frac{1}{\alpha}\frac{1}{t^{1/\alpha}}\frac{t^{1/\alpha}}{|x|}H_{2,2}^{1,1}\left[\frac{|x|}{Kt^{1/\alpha}}\big{|}_{(1,1),(1,\frac{1}{2})}^{(1,\frac{1}{\alpha}),(1,\frac{1}{2})}\right]$
$=\frac{1}{\alpha|x|}\frac{1}{2\pi i}\int_{c-i\infty}^{c+i\infty}\frac{\Gamma(1+\frac{s}{\alpha})\Gamma(-\frac{s}{2})}{\Gamma(-s)\Gamma(1+\frac{s}{2})}\left(\frac{|x|}{Kt^{1/\alpha}}\right)^{-s}{\rm d}s$
where $c$ in the contour is such that $-\alpha<c<0$. Replacing $\frac{s}{\alpha}$ by $s$, observing that $\alpha$ coming from $s~{}{\rm d}s$ is canceled with $\alpha$ sitting outside, we have the following:
$p(x,t)=\frac{1}{|x|}\frac{1}{2\pi i}\int_{c^{\prime}-i\infty}^{c^{\prime}+i\infty}\frac{\Gamma(1+s)\Gamma(-\frac{\alpha}{2}s)}{\Gamma(-\alpha s)\Gamma(1+\frac{\alpha}{2}s)}\left(\frac{|x|}{Kt^{1/\alpha}}\right)^{-\alpha s}{\rm d}s,-1<c^{\prime}<0.$
By using the duplication formula for gamma functions, we have
$\Gamma(-\alpha s)=\Gamma(2(-\frac{\alpha}{2}s))=2^{-\alpha s-1}\pi^{-\frac{1}{2}}\Gamma(-\frac{\alpha}{2}s)\Gamma(\frac{1}{2}-\frac{\alpha}{2}s)$
so that one $\Gamma(-\frac{\alpha}{2}s)$ is canceled. Then,
$p(x,t)=\frac{2\pi^{\frac{1}{2}}}{|x|}\frac{1}{2\pi i}\int_{c^{\prime}-i\infty}^{c^{\prime}+i\infty}\frac{\Gamma(1+s)}{\Gamma(\frac{1}{2}-\frac{\alpha}{2}s)\Gamma(1+\frac{\alpha}{2}s)}\left(\frac{|x|}{2Kt^{1/\alpha}}\right)^{-\alpha s}{\rm d}s.$
Evaluating the H-function as the sum of the residues at the poles of $\Gamma(1+s)$, which are at $s=-1-\nu,\nu=0,1,....$, we have the following series:
$p(x,t)$ $=\frac{2\pi^{\frac{1}{2}}}{|x|}(\frac{|x|^{\alpha}}{(2K)^{\alpha}t})$
×$\times\sum_{\nu=0}^{\infty}\frac{(-1)^{\nu}}{\nu!}\frac{1}{\Gamma(\frac{1+\alpha}{2}+\frac{\alpha}{2}\nu)\Gamma(1-\frac{\alpha}{2}(1+\nu))}\left(\frac{|x|^{\alpha}}{(2K)^{\alpha}t}\right)^{\nu},\alpha\neq 1,\alpha\neq 2.$
4 Discussion
The first solar neutrino experiment led by Raymond Davis Jr. showed a deficit of neutrinos relative to the solar model prediction, referred to as the solar neutrino problem since the 1970s. The Kamiokande experiment led by Masatoshi Koshiba successfully observed solar neutrinos, as first reported in 1980s. The solar neutrino problem was solved due to neutrino oscillations by comparing the Super-Kamiokande and Sudbury Neutrino Observatory results. While recent decades have offered tremendous advances in solar neutrinos across the fields of solar physics (Buldgen et al., 2024), nuclear physics (Bertulani et al., 2022), and neutrino physics (Slad, 2023), many lingering mysteries remain.
This paper takes advantage of publicly available solar neutrino Super-Kamiokande data and analyses them by applying diffusion entropy analysis and standard deviation analysis. The result is a scaling exponent $\delta<1$ indicating anomalous diffusion of solar neutrinos in terms of Lévy flights. Based on this result the paper developes the probability density function for neutrino flights and derives the respective differential equation in terms of a fractional Fokker-Planck equation. Accordingly, the closed form analytic representation of the neutrino power density function is given as a Fox H-function that can be used for further numerical exercises for the benefit of solar neutrino physics.
Acknowledgements
The authors are grateful for the support in the diffusion entropy analysis and standard deviation analysis by Dr. Alexander Haubold while he was doing his research at the Department of Computer Science, Columbia University, New York (USA). The results of diffusion entropy analysis were independently confirmed by the research of Dr. Nicy Sebastian, Department of Statistics, St Thomas College, Thrissur, University of Calicut, Kerala (India).
Figure 3: The Standard Deviation Analysis (SDA) of the ${}^{8}B$ solar neutrino data from the SuperKamiokande I and II experiment.
Figure 4: The Diffusion Entropy Analysis (DEA) of the ${}^{8}B$ solar neutrino data from the SuperKamiokande I and II experiment.
Figure 5: The Standard Deviation Analysis (SDA) of the $hep$ solar neutrino data from the SuperKamiokande I and II experiment.
Figure 6: The Diffusion Entropy Analysis (DEA) of the $hep$ solar neutrino data from the SuperKamiokande I and II experiment.